A study was designed to compare the attitudes of two groups of nursing students towards computers. Group 1 had previously taken a statistical methods course that involved significant computer interaction. Group 2 had taken a statistic methods course that did not use computers. The students' attitudes were measured by administering the Computer Anxiety Rating Scale (CARS). A random sample of 11 nursing students from Group 1 resulted in a mean score of 49.8 with a standard deviation of 5.7. A random sample of 19 nursing students from Group 2 resulted in a mean score of 55.2 with a standard deviation of 2.2. Can you conclude that the mean score for Group 1 is significantly lower than the mean score for Group 2? Let μ1 represent the mean score for Group 1 and μ2 represent the mean score for Group 2. Use a significance level of α=0.01 for the test. Assume that the population variances are equal and that the two populations are normally distributed.

Step 2 of 4 : Compute the value of the t test statistic. Round your answer to three decimal places.

Step 3 of 4 : Determine the decision rule for rejecting the null hypothesis H0. Round your answer to three decimal places.

Step 4 of 4: State the test's conclusion.
Answer #1

H0 : mu1 = mu2
Ha: mu1 < mu2

2)

test statsitics:

Sp = sqrt(((n1-1)*s1^2+(n2-1)*s2^2)/(n1+n2-2))
= sqrt(((11- 1) 8 5.7^2 + (19-1)*2.2^2)/(11+19-2))
= 3.836


t = (x1 -x2)/sp *sqrt(1/n1+1/n2)
= (49.8 - 55.2)/(3.836 *sqrt(1/11 + 1/19))
= -3.716

3)

Reject H0 if t < -2.467

4)

Reject H0